``Before, solving the problem, I just want to give you the heads up. It's
Systems of Equation: Application. I can't find the specific title to relate
with my story problems. My chapter in the book is Chapter Eight, Solving Systems by Substiution, Additional. Hope it helps!
*Two families attend the town's ice cream social every summer. This year one family piad $9.75 for 3 adult tickets and 5 children's tickets. The other family paid $10.75 for 5 adult tickets and 1 child's ticket. Find the price of a child's ticket.*

Answer by Mathtut(3670)

let c and a be the prices childrens and adult tickets respectively
:
5c+3a=9.75............eq 1
1c+5a=10.75...........eq 2
:
rewrite eq 2 to c=10.75-5a and plug this value into eq 1
:
5(10.75-5a)+3a=9.75
:
53.75-25a+3a=9.75
:
-22a=-44
:
a=$2
:
c=10.75-5(2)=$.75